OpenStudy (anonymous):

A stuntman falls from a hot-air balloon that is hovering at a constant altitude, 100 ft above a lake. A television camera on shore, 200 ft from a point directly below the balloon, follows his descent. At what rate is the angle of elevation, theta, of the camera changing 2 seconds after the stuntman falls? (Neglect the height of the camera.)

OpenStudy (anonymous):

\[ \tan(\theta) = y/x \] Implicitly differentiate with respect to \(t\). What equation do you get?

OpenStudy (anonymous):

This?\[\sec ^{2}\theta \frac{ d \theta }{ dt } =\frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } }\]

OpenStudy (anonymous):

You'd have to use the quotient rule actually.

OpenStudy (anonymous):

\[ (y/x)' = \frac{y'x-yx'}{x^2} \]

OpenStudy (anonymous):

so then...\[\sec ^{2}\theta \frac{ d \theta }{ dt }=\frac{ \frac{ dy }{ dt }x-y \frac{ dx }{ dt } }{ x ^{2} }\]

OpenStudy (anonymous):

Yeah, now since \(x\) is constant with respect to time, we know \(x'=0\).

OpenStudy (anonymous):

Also, since \(x = \cos\theta\) we know \(\sec^2\theta = x^{-2}\).

OpenStudy (anonymous):

This will simplify things quite a bit.

OpenStudy (anonymous):

why is x=cosθ?

OpenStudy (anonymous):

Ooops. I meant to say the \(z\cos\theta =x\) where \(z\) is the hypotenuse.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

So \(\sec^2\theta = z^2/x^2\)

OpenStudy (anonymous):

\( z^2/x^2 =( y^2+x^2)/x^2\)

OpenStudy (anonymous):

What equation are you ending up with for \(\theta'\)?

OpenStudy (anonymous):

\[\frac{ d \theta }{ dt }=\frac{ \frac{ dy }{ dt }x }{ (y ^{2}+x ^{2}) }\]

OpenStudy (anonymous):

Yeah. Now you are given y and x, all you need is y'

OpenStudy (anonymous):

the gravitational constant -32.174ft/sec^2?

OpenStudy (anonymous):

That's acceleration.

OpenStudy (anonymous):

You want speed.

OpenStudy (anonymous):

do you mean velocity such as taking the integral of 32.174ft/sec^2 and coming up with -32.174t ft/sec

OpenStudy (anonymous):

\[ y' = \int_0^2y''dt \]In this case, \(y''\) is the acceleration due to gravity.

OpenStudy (anonymous):

so 64.384

OpenStudy (anonymous):

Yeah. Now you can solve for \(\theta'\)

OpenStudy (anonymous):

I mean you have everything you need now, since: \[ \theta' = \frac{y'x}{x^2+y^2} \]

OpenStudy (anonymous):

so...\[\frac{ d \theta }{ dt }=\frac{ 12869.6 }{ 50000 } = 0.257392\]

OpenStudy (anonymous):

or no I forgot the negative sign y'=-64.348 so... it should be -0.257392 correct?

OpenStudy (anonymous):

It's in radians, I believe.

OpenStudy (anonymous):

yes rads

OpenStudy (anonymous):

but otherwise the answer is: The rate of change in the angle of the camera at 2 seconds after the stuntman jumps is -0.257392rads correct?

OpenStudy (anonymous):

You could try differentiating: \[ z\sin\theta = y \]to see if the answer is consistent.

OpenStudy (anonymous):

actually that would be a bit complicated.

OpenStudy (anonymous):

yes it seems like it would be :-)

OpenStudy (anonymous):

Wait, actually there might be a problem with the answer you got.

OpenStudy (anonymous):

We should probably use \(y(t=2)\) rather than \(y(t=0)\).

OpenStudy (anonymous):

Why is that?

OpenStudy (anonymous):

Because \(y=y(t)\) and we overlooked that fact.

OpenStudy (anonymous):

It's a function of \(t\) and we set \(t=2\).

OpenStudy (anonymous):

\[ \int_0^2y'dt = y(2)-y(0)\implies y(2) = \int_0^2y'dt + y(0) \]

OpenStudy (anonymous):

This time though the integrand (\(y'\) ) is not constant. \[ y'(t) = y''\cdot t= -32.174t \]

OpenStudy (anonymous):

Sorry I didn't notice that at first.

OpenStudy (anonymous):

how is y'(t)=y'' * t =-32.174t?

OpenStudy (anonymous):

it it because V=gt? where g=gravitational acceleration and t=time?

OpenStudy (anonymous):

Okay so basically....\[ \int_0^u y''dt = y'(u) - y'(0) \]By the fundamental theorem of calculus (\(u\) is just a dummy variable). At the start, the velocity is 0, so \[ y(u) = \int_0^u y''dt = -32.174\int_0^udt =-32.174 \left[t \bigg|_0^u \right] = -32.174u \]

OpenStudy (anonymous):

Replacing \(t\) back into \(u\): \[ y'(t) = -32.174t \]

OpenStudy (anonymous):

Likewise: \[ y(2) -y(0)= \int_0^2 y'' dt \implies y(2) - 100 = \int^2_0 -32.174tdt \]

OpenStudy (anonymous):

The whole \(v=gt\) comes from calculus.

OpenStudy (anonymous):

The derived formula are fine if you remember the preconditions for them, otherwise just use the definitions to get things done.

OpenStudy (anonymous):

In math section, I avoid the derived formula and use mainly definitions / laws / principles.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so it should be:\[\frac{ d \theta }{ dt }=\frac{ -32.174t*200 }{10000+40000 }\]

OpenStudy (anonymous):

No, you never solved for \(y(2)\)

OpenStudy (anonymous):

\[ \frac{d\theta}{dt}(2) = \frac{\frac{dy}{dt}(2)\cdot x(2)}{[y(2)]^2+[x(2)]^2} \]

OpenStudy (anonymous):

y(2)=-16.087t^2

OpenStudy (anonymous):

You are evaluating when \(t=2\), there should be no \(t\)s left.

OpenStudy (anonymous):

A definite integral always returns a number that is constant

OpenStudy (anonymous):

so y(2)=-64.384

OpenStudy (anonymous):

oh -64.348

OpenStudy (anonymous):

You're saying they're underground? That makes no sense.

OpenStudy (anonymous):

\[ y(2) -y(0)= \int_0^2 y'' dt \implies y(2) - 100 = \int^2_0 -32.174tdt \]

OpenStudy (anonymous):

\[ y(2) = 100 + -32.174\left[ \frac{t^2}{2}\bigg|_0^2\right] \]

OpenStudy (anonymous):

if y'=y''*t or -32.174t the integral of y' with limits of integration of [0,2] is -16.087{(2)^2}or -64.348

OpenStudy (anonymous):

Yeah but you are forgetting that \(y(0) \neq 0 \)

OpenStudy (anonymous):

it is = 100

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

so y(2) = 35.652

OpenStudy (anonymous):

Now that makes sense.

OpenStudy (anonymous):

so....\[\frac{ d \theta }{ dt }=\frac{ -64.348(200) }{ (35.653)^{2}+40000 }\]

OpenStudy (anonymous):

so then θ' is -0.3118305217588708 or in answer form decreasing at a rate of 0.3118305217588708rads

OpenStudy (anonymous):

Yeah that looks right.

OpenStudy (anonymous):

okay, now I'm going to try and figure out were I went wrong the first time I tried to solve this. Thank you so much for your help :-)